OpenStudy (anonymous):

MEDAL $ U _A ball is thrown into the air with an initial upward velocity of 60 ft/s. Its height(H) in feet after t seconds is given by the function h=-16t2+60t+6.What will the height be at t=3 seconds? 1.35 ft b.40ft c.42ft d. 45ft

OpenStudy (shiraz14):

Sub t=3 in the equation, we have: h=-16(3^2)+(60x3)+6 = -144+180+6 = 42ft So the answer should be C.

OpenStudy (anonymous):

I have one more , could you help ?

OpenStudy (anonymous):

One more rectangular shaped piece of metal siding needs to be cut to cover the exterior of a pole barn. The area of the piece is 30 FT 2. The lenght is 1 less than 3 times the width.How wide should the medal piece be ? Round to the nearest hundredth of a foot . a. 3.33 ft b.4.3 ft c.1 ft d.30 ft

OpenStudy (anonymous):

a?

OpenStudy (shiraz14):

Let the width be x. Then the length will be 3x -1. So the area would be x(3x-1) = 3x^2-x=30 ft^2. We thus have the equation 3x^2 - x - 30 = 0. Using the quadratic formula, we would have x = 3.33 ft. The correct answer is thus A.

OpenStudy (anonymous):

YEs I got it right ! Ty

OpenStudy (shiraz14):

You're welcome

OpenStudy (anonymous):

Your a def pro. If you ever need help in any subjects outside of math I could help u too. I get straight a's

OpenStudy (anonymous):

I'm posting 5 more examples on the math Board if you would like to catch some more medals